# CausalDynamics: A large-scale benchmark for structural discovery of dynamical causal models

### v1

@@ 4.6 Limitations
PCMCI+ F-PCMCI VARLiNGAM DYNOTEARS NGC TSCI CUTS+
@@
MAOOAM .69  /  .88 .50  /  .81 .50  /  .81 .62  /  .86 .49  /  .81 .58  /  .84 .50  /  .81
ENSO .50  /  .81 .51  /  .81 .51  /  .81 .50  /  .81 .50  /  .81 .50  /  .81 .49  /  .81
@@ 5 Experiments
To showcase how SOTA algorithms handle the various challenges of our benchmark, we evaluate the approaches described in Section 3.2, reporting a selection of the results in Table 2 and the full results online at https://huggingface.co/datasets/kausable/CausalDynamics. The default setup for the simple case refers to ODE systems ($\delta=0$) with no unobserved confounder. In the coupled systems tier, the default setup refers to dynamics with $n=10$ coupled ODEs ($\delta=0$), no confounder, no time-lag ($\tau=0$), and no standardization. We performed an initial hyperparameter search for most of the baseline algorithms and provide the results in Section E.1.
We find methods that do not require much fine tuning, such as the non-DL-based algorithms, perform better as illustrated in Table 2 and Figure 4, and observed by [18]. This is striking for the coupled atmosphere-ocean model, where each node represents a 2-dimensional spatial field. Methods that only consider 1-dimensional time-series, perform best, compared to DL-based methods, which claim to have a notion of space. We also find some evidence that topology-based methods, such as TSCI, perform better than purely DL-based approaches, for confounded and higher dimensional systems.
Overall, existing algorithms show shortcomings when confronted with coupled dynamical and physical systems as shown in Figure 5 for a random single graph realization. Across methods we find that autocorrelation is inferred where none exists. The baselines also appear conservative, predicting dense adjacency in the presence of nonstationary dynamics. In the climate case, methods perform well but do not recover the full coupling between ocean basins. Nevertheless, scores are the highest for the climate examples, possibly due to the noise in the generated data which might benefit some methods in recovering the underlying graph structure. Surprisingly in Figure 4, we observe minimal improvement in the standardized case, especially for methods like VARLiNGAM that exploit topological ordering.
Finally, we note that for sparse graphs, AUROC can be deceptively high because the false-positive rate remains low due to a higher number of true-negatives. Thus, for coupled systems AUPRC scores might be more representative of the true performance due to the penalization of false positives. Other metrics like Structural Hamming Distance [91] or Structural Intervention Distance [92], suffer from different limitations as they generally only return absolute error counts independent from the edge density or graph size, thus making a comparison across graph hierarchies difficult.
@@ 6 Conclusion
We present CausalDynamics, an extensible data generation framework that we use to construct the largest benchmark dataset with over 5300 preprocessed graphs of increasing complexity. We structure our benchmark as a tiered system ranging from simple three-dimensional dynamical systems to pseudo-real physical systems. We provide a plug-and-play workflow to facilitate the development of novel causal discovery methods across various domains. Evaluating a set of state-of-the-art causal discovery algorithms on CausalDynamics shows that many advanced DL-based algorithms are outperformed by simpler methods, notably on high-dimensional datasets, highlighting a need for future method development. We believe our work provides the necessary foundation for the advancement of causal discovery algorithms that are applicable in high-dimensional, nonlinear and dynamical settings.
Software and Data
Acknowledgments
We would like to thank Lars Lorch and Johannes Veith for discussions and their valuable insights. The authors acknowledge funding, computing, and storage resources from the NSF Science and Technology Center (STC) Learning the Earth with Artificial Intelligence and Physics (LEAP) (Award #2019625) and Department of Energy (DOE) Advanced Scientific Computing Research (ASCR) program (DE-SC0022255). J.B., and P.G. acknowledge funding from the Zegar Family Foundation. J.N. also acknowledges funding from Columbia-Dream Sports AI Innovation Center. Finally, this work was supported by the EXIST-Gründerstipendium by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) in collaboration with the SRH Gründerinstitut.
@@ Appendix A Getting started
All results, code examples and descriptions rely on version 0.1.1 of the CausalDynamics Python package.
@@ A.3 Generate data
init_ratios = [1, 1] # Set ratios of dynamical systems
# and periodic drivers at root nodes.
@@ B.1 System initialization
If the dynamical system drivers are selected at random, the algorithm chooses randomly from the set of available 3D systems (see Algorithm 3). Users can also choose whether to include periodic functions as drivers by initializing r appropriately. In this case, sinusoidal trajectories are generated as outlined in Algorithm 4. The generated data of the system’s drivers is randomly assigned to the respective root nodes, and all non-root nodes initial states are set to zero (see Algorithm 5).
@@
1:$T$ (num_timesteps), $N$ (num_nodes), $r$ (init_ratios), name (system_name), $D$ (node_dim), $l$ (time_lag)
2:$\text{init}\in\mathbb{R}^{T\times N\times D}$ (initial values tensor)
@@
4:     $T_{ext}\leftarrow T+l$triangleright Extended time for lag calculation
@@
8:$n_{sys},n_{sin}\leftarrow\text{allocate\_elements\_based\_on\_ratios}(N,r)$triangleright Divide nodes by type
@@
10:     $d_{sys}\leftarrow\text{solve\_random\_systems}(T_{ext},n_{sys})$triangleright Generate random system trajectories
@@
[…]
@@
14:$d_{sin}\leftarrow\text{drive\_sin}(T_{ext},n_{sin},D)$triangleright Generate sinusoidal trajectories
15:$\text{init}\leftarrow\text{concat}(d_{sys},d_{sin},\text{dim}=1)$triangleright Combine trajectories
16:$\text{init}\leftarrow\text{init}[:,\text{randperm}(N),:]$triangleright Randomly permute nodes
17:if $l>0$ then
@@
21:end if
22:return init
@@
1:$T$ (num_timesteps), $N$ (num_nodes)
2:$\text{data}\in\mathbb{R}^{T\times N\times D}$ (trajectory tensor)
[…]
5:$\text{data}\leftarrow[]$triangleright Initialize empty trajectory list
6:for $i\leftarrow 0$ to $N-1$ do
[…]
8:     $\text{counter}\leftarrow 0$
@@
17:     $\text{data}.\text{append}(\text{sol})$triangleright Add trajectory to list
@@
$\text{data}\leftarrow\text{data}.\text{permute}(1,0,2)$triangleright Reshape to [T, N, D]
@@
1:$T$ (num_timesteps), $N$ (num_nodes), $D$ (node_dim), $P_{max}$ (max_num_periods)
2:$\text{data}\in\mathbb{R}^{T\times N\times D}$ (sinusoidal trajectory tensor)
3:$\text{amplitude}\leftarrow 2\times\text{rand}(N,D)-1$triangleright Random amplitudes in $[-1,1]$
[…]
5:$\text{data}\leftarrow\text{zeros}(T,N,D)$triangleright Initialize output tensor
6:if $N>0$ then
@@
1:$\text{init}\in\mathbb{R}^{T\times N\times D}$ (initial values tensor), $A\in\mathbb{R}^{N\times N}$ or $\mathbb{R}^{2N\times N}$ (adjacency matrix), standardize (boolean)
2:$x\in\mathbb{R}^{T\times N\times D}$ (initialized tensor with values for root nodes and zero else)
[…]
7$\triangleright$ With time lag edges
8:     $A_{now}\leftarrow A[:A.\text{shape}[0]/2]$triangleright Extract current time step connections
9:     $A_{past}\leftarrow A[A.\text{shape}[0]/2:]$triangleright Extract past time step connections
@@
16:     $\mu\leftarrow\text{mean}(x,\text{dim}=0)$triangleright Compute mean over time
17:     $\sigma^{2}\leftarrow\text{var}(x,\text{dim}=0)$triangleright Compute variance over time
18:     $x\leftarrow\frac{x-\mu}{\sqrt{\sigma^{2}}}$triangleright Standardize results
@@ B.2 MLP propagation
1:$A\in\{0,1\}^{n\times n}$ (adjacency matrix), $W\in\mathbb{R}^{n\times d\times d}$ (weight tensor), $b\in\mathbb{R}^{n\times d}$ (bias tensor), $\text{init}\in\mathbb{R}^{T\times n\times d}$ (initial values), $\text{standardize}\in\{\text{True},\text{False}\}$
2:$x\in\mathbb{R}^{T\times n\times d}$ (propagated values)
3:$T,n,d\leftarrow\text{shape}(\text{init})$triangleright Get dimensions
4:$x\leftarrow\text{initialize\_x}(\text{init},A)$triangleright Initialize state tensor, see Algorithm 5
5:$A\leftarrow A^{T}$triangleright Transpose adjacency matrix for easier indexing
6:for $i\leftarrow n-1$ to 00 do $\triangleright$ Process nodes in reverse topological order
@@
10:     $m_{i}\leftarrow A[i].\text{bool}()$triangleright Get incoming edge mask
@@
12:         $W_{sel}\leftarrow W[m_{i}]$triangleright Select weights for incoming edges
13:         $b_{sel}\leftarrow b[m_{i}]$triangleright Select biases for incoming edges
14:         $x_{sel}\leftarrow x[:,m_{i}]$triangleright Select input values
15:         $y\leftarrow\text{matmul}(W_{sel},x_{sel})+b_{sel}$triangleright MLP transformation
16:         $y\leftarrow\text{sum}(y,\text{dim}=1)$triangleright Aggregate incoming signals
17:         if standardize then
18:              $\mu\leftarrow\text{mean}(y,\text{dim}=0)$triangleright Compute mean over time
19:              $\sigma^{2}\leftarrow\text{var}(y,\text{dim}=0)$triangleright Compute variance over time
20:              $y\leftarrow\frac{y-\mu}{\sqrt{\sigma^{2}}}$triangleright Standardize results
21:         end if
22:         $x[:,i]\leftarrow x[:,i]+y$triangleright Update node values
23:     end if
24:end for
25:return $x$
@@ C.1 XRO: eXtended nonlinear Recharge Oscillator model
The XRO model is a Python implementation [83] of a recharge oscillator (RO) model for El Niño-Southern Oscillation (ENSO) [85] coupled to stochastic-deterministic models for other climate modes (M) (as shown in Figure 11) which allows for a two-way interaction. A detailed description of the model is given by [86]. In the following we will provide an overview.
@@
The qgs***https://github.com/Climdyn/qgs/ library provides a Python implementation of a simplified, first-order coupled atmosphere–ocean model, based in part on the Modular Arbitrary-Order Ocean-Atmosphere Model (MAOOAM) [87]. It represents a two-layer quasi-geostrophic atmosphere ($\psi^{1},\psi^{3}$), represented by barotropic $\psi_{a}$ and baroclinic $\theta_{a}$ streamfunctions, mechanically and thermally coupled to a shallow-water ocean component ($\psi_{o}$), with interactions driven by wind forcing alongside radiative and heat exchanges (see Figure 12a).
@@
Here, $W,K,d$ and $s$ denote the coupling coefficients governing ocean–atmosphere interactions; $a,g,b$ and $c$ are the inner-product coefficients for the atmospheric Fourier modes and $M,O,C,N$ those for the ocean. A detailed model description is given in [87] and https://qgs.readthedocs.io/en/latest/files/model/maooam_model.html.

### v2

@@ 4.6 Limitations
PCMCI+ F-PCMCI VARLiNGAM DYNOTEARS NGC TSCI CUTS+ RCD GRaSP TCDF
@@
[…]
Table 3: Baseline SHD ($\downarrow$) score across different experiments in the hierarchy of increasingly complex dynamical systems. The scores are averaged over all generated graphs within a specific experiment. The experiments reported here consider one causal problem at a time, while keeping other factors at the default setting. The full results, including all possible experimental permutations, are found in our data repository.
PCMCI+ F-PCMCI VARLiNGAM DYNOTEARS NGC TSCI CUTS+ RCD GRaSP TCDF
Experiments Simple
[…]
Coupled
[…]
Climate
[…]
@@ 5 Experiments
To showcase how SOTA algorithms handle the various challenges of our benchmark, we evaluate the approaches described in Section 3.2, reporting a selection of the results in Tables 2-3 and the full results online at https://huggingface.co/datasets/kausable/CausalDynamics. The default setup for the simple case refers to ODE systems ($\delta=0$) with no unobserved confounder. In the coupled systems tier, the default setup refers to dynamics with $n=10$ coupled ODEs ($\delta=0$), no confounder, no time-lag ($\tau=0$), and no standardization. We performed an initial hyperparameter search for most of the baseline algorithms and provide the results in Section E.1.
We find methods that do not require much fine tuning, such as the non-DL-based algorithms, perform better as illustrated in Tables 2-3 and Figure 4, and observed by [20]. This is striking for the coupled atmosphere-ocean model, where each node represents a 2-dimensional spatial field. Methods that only consider 1-dimensional time-series, perform best, compared to DL-based methods, which claim to have a notion of space. We also find some evidence that topology-based methods, such as TSCI, perform better than purely DL-based approaches, for confounded and higher dimensional systems.
Overall, existing algorithms show shortcomings when confronted with coupled dynamical and physical systems as shown in Figure 5 for a random single graph realization. Across methods we find that autocorrelation is inferred where none exists. The baselines also appear conservative, predicting dense adjacency in the presence of nonstationary dynamics. In the climate case, methods perform well but do not recover the full coupling between ocean basins. Nevertheless, scores are the highest for the climate examples, possibly due to the noise in the generated data which might benefit some methods in recovering the underlying graph structure. Surprisingly in Figure 4, we observe minimal improvement in the standardized case, especially for methods like VARLiNGAM that exploit topological ordering. We performed additional ablation studies that consider the effect of time subsampling, partial observability, and varying edge-level activation function in Appendix F.
Finally, we note that for sparse graphs, AUROC can be deceptively high because the false-positive rate remains low due to a higher number of true-negatives. Thus, for coupled systems AUPRC scores might be more representative of the true performance due to the penalization of false positives. Other metrics like SHD [97, 98] or Structural Intervention Distance [99], suffer from different limitations as they generally only return absolute error counts independent from the edge density or graph size, thus making a comparison across graph hierarchies difficult.
@@ 6 Conclusion
We present CausalDynamics, an extensible data generation framework that we use to construct the largest benchmark dataset with over 14000 preprocessed graphs of increasing complexity. We structure our benchmark as a tiered system ranging from simple three-dimensional dynamical systems to pseudo-real physical systems. We provide a plug-and-play workflow to facilitate the development of novel causal discovery methods across various domains. Evaluating a set of state-of-the-art causal discovery algorithms on CausalDynamics shows that many advanced DL-based algorithms are outperformed by simpler methods, notably on high-dimensional datasets, highlighting a need for future method development. We believe our work provides the necessary foundation for the advancement of causal discovery algorithms that are applicable in high-dimensional, nonlinear and dynamical settings.
Acknowledgment
We would like to thank Lars Lorch and Johannes Veith for discussions and their valuable insights. The authors acknowledge funding, computing, and storage resources from the NSF Science and Technology Center (STC) Learning the Earth with Artificial Intelligence and Physics (LEAP) (Award #2019625), and Department of Energy (DOE) Advanced Scientific Computing Research (ASCR) program (DE-SC0022255). JN also acknowledges funding from Columbia-Dream Sports AI Innovation Center. The project on which this report is based was funded by the Federal Ministry for Economic Affairs and Climate Action under the funding code 03EGTBW076. The responsibility for the content of this publication lies with the authors. Finally, the project was funded by the BW Pre-Seed program of the State of Baden-Württemberg, funded by the Ministry of Economics, Labour and Tourism and the L-Bank.
@@ Appendix A Getting started
All results, code examples and descriptions rely on version 1.0.0 of the CausalDynamics Python package.
@@ A.3 Generate data
init_ratios = [1, 1, 1]# Set ratios of dynamical systems,
# periodic, and linear drivers at root nodes.
@@
activations_names = [’identity’, ’sin’, ’sigmoid’, ’tanh’, ’relu’]
# Activation names sampled uniformly per edge
@@ A.5 Baseline evaluation
We also provide the complete metrics for each graph in our HuggingFace data repository https://huggingface.co/datasets/kausable/CausalDynamics, but analyzing each of the 14k+ graphs can be too granular. As a result, we have added a built-in, easy-to-use diagnosis script that allows users to quickly analyze results on the experiment-level. For instance, if users are interested in understanding the effects of unobserved confounders (confounders=True) on the performance of different causal discovery algorithms, especially in light of high internal noise/stochasticity (noise=2.0), they can run the following:
# Diagnostic tool to evaluate each experiment
python diagnose.py --exp_dir data/simple/noise=2.00_confounder=True
@@ B.1 System initialization
If the dynamical system drivers are selected at random, the algorithm chooses randomly from the set of available 3D systems (see Algorithm 3). Users can also choose whether to include periodic or linear functions as drivers by initializing r appropriately. In this case, sinusoidal or linear trajectories are generated as outlined in Algorithm 4 and Algorithm 5. The generated data of the system’s drivers is randomly assigned to the respective root nodes, and all non-root nodes initial states are set to zero (see Algorithm 6).
@@
1:$T$ (num_timesteps), NN (num_nodes), $r$ (init_ratios), name (system_name), $D$ (node_dim), $l$ (time_lag)
[…]
@@
4:  $T_{ext}\leftarrow T+l$ $\triangleright$ Extended time for lag calculation
@@
9:if $\text{system\_name}=\text{"random"}$ then
10:  $d_{sys}\leftarrow\text{solve\_random\_systems}(T_{ext},n_{sys})$ $\triangleright$ Generate random system trajectories
@@
12:  $d_{sys}\leftarrow\text{solve\_system}(T_{ext},n_{sys},\text{system\_name})$ $\triangleright$ Generate named system trajectories
@@
[…]
18:if $l>0$ then
19:  $\text{init\_now}\leftarrow\text{init}[:T,:,:]$ $\triangleright$ Current time steps
20:  $\text{init\_future}\leftarrow\text{init}[l:T_{ext},:,:]$ $\triangleright$ Future time steps for lagged edges
21:  $\text{init}\leftarrow\text{concat}(\text{init\_now},\text{init\_future},\text{dim}=1)$ $\triangleright$ Combine current and future
22:end if
23:return init
@@
1:$T$ (num_timesteps), NN (num_nodes)
@@
[…]
6:for $i\leftarrow 0$ to N1N-1 do
7:  $\text{sol}\leftarrow\text{None}$
8:  $\text{counter}\leftarrow 0$
9:  while $\text{sol}=\text{None}$ and $\text{counter}<\text{max\_retry}$ do $\triangleright$ Handle failing integration
10:   $\text{system}\leftarrow\text{random\_choice}(\text{names}\setminus\{\text{s}[i\bmod|\text{s}|]\})$
11:   $\text{sol}\leftarrow\text{solve\_single\_system}(\text{system},T)$ $\triangleright$ Solve system ODEs
@@
17:  $\text{data}.\text{append}(\text{sol})$ $\triangleright$ Add trajectory to list
@@
[…]
@@
1:$T$ (num_timesteps), NN (num_nodes), $D$ (node_dim), $P_{max}$ (max_num_periods)
@@
[…]
6:if $N>0$ then
7:  $\text{max\_time}\leftarrow P_{max}\times 2\pi\times\text{rand}(N,D)$ $\triangleright$ Random max times
8:  $\text{time}\leftarrow\text{linspace}(0,\text{max\_time},T)$ $\triangleright$ Generate time points
9:  for $i\leftarrow 0$ to T1T-1 do
10:   $\text{data}[i,:,:]\leftarrow\text{amplitude}\times\sin(\text{time}[i]+\text{phase\_shift})$ $\triangleright$ Compute sine values
11:  end for
12:end if
13:return data
Algorithm 5 Generate Linear Driver Trajectories
1:$T$ (num_timesteps), NN (num_nodes), $D$ (node_dim), $m_{\min},m_{\max}$ (slope range), $b_{\min},b_{\max}$ (intercept range)
[…]
6:if $N>0$ then
7:  $\text{max\_time}\leftarrow\text{rand}(N,D)$ $\triangleright$ Random max times in $[0,1]$
8:  $\text{time}\leftarrow\text{linspace}(0,\text{max\_time},T)$ $\triangleright$ Generate time points
9:  for $i\leftarrow 0$ to T1T-1 do
10:   $\text{data}[i,:,:]\leftarrow m\times\text{time}[i]+b$ $\triangleright$ Compute linear values
@@
5:  $\text{root\_nodes}\leftarrow\text{get\_root\_nodes\_mask}(A)$ $\triangleright$ Identify nodes with no incoming edges
6:  $x[:,\text{root\_nodes},:]\leftarrow\text{init}[:,\text{root\_nodes},:]$ $\triangleright$ Set values only for root nodes
@@
8:  $A_{now}\leftarrow A[:A.\text{shape}[0]/2]$ $\triangleright$ Extract current time step connections
9:  $A_{past}\leftarrow A[A.\text{shape}[0]/2:]$ $\triangleright$ Extract past time step connections
10:  $\text{root\_now}\leftarrow\text{get\_root\_nodes\_mask}(A_{now})$ $\triangleright$ Current time root nodes
11:  $\text{root\_past}\leftarrow\text{get\_root\_nodes\_mask}(A_{past})$ $\triangleright$ Past time root nodes
12:  $\text{root\_nodes}\leftarrow\text{concat}(\text{root\_now},\text{root\_past},\text{dim}=0)$ $\triangleright$ Combine masks
13:  $x[:,\text{root\_nodes},:]\leftarrow\text{init}[:,\text{root\_nodes},:]$ $\triangleright$ Set values only for root nodes
@@
16:  $\mu\leftarrow\text{mean}(x,\text{dim}=0)$ $\triangleright$ Compute mean over time
17:  $\sigma^{2}\leftarrow\text{var}(x,\text{dim}=0)$ $\triangleright$ Compute variance over time
18:  $x\leftarrow\frac{x-\mu}{\sqrt{\sigma^{2}}}$ $\triangleright$ Standardize results
@@ B.2 MLP propagation
[…]
6:for $i\leftarrow n-1$ to 0 do $\triangleright$ Process nodes in reverse topological order
@@
10:  $m_{i}\leftarrow A[i].\text{bool}()$ $\triangleright$ Get incoming edge mask
@@
12:   $W_{sel}\leftarrow W[m_{i}]$ $\triangleright$ Select weights for incoming edges
13:   $b_{sel}\leftarrow b[m_{i}]$ $\triangleright$ Select biases for incoming edges
14:   $x_{sel}\leftarrow x[:,m_{i}]$ $\triangleright$ Select input values
15:   $y\leftarrow\text{matmul}(W_{sel},x_{sel})+b_{sel}$ $\triangleright$ MLP transformation
16:   $y\leftarrow\phi[i](y)$ $\triangleright$ Apply activation (optional)
17:   $y\leftarrow\text{sum}(y,\text{dim}=1)$ $\triangleright$ Aggregate incoming signals
18:   if standardize then
19:     $\mu\leftarrow\text{mean}(y,\text{dim}=0)$ $\triangleright$ Compute mean over time
20:     $\sigma^{2}\leftarrow\text{var}(y,\text{dim}=0)$ $\triangleright$ Compute variance over time
21:     $y\leftarrow\frac{y-\mu}{\sqrt{\sigma^{2}}}$ $\triangleright$ Standardize results
22:   end if
23:   $x[:,i]\leftarrow x[:,i]+y$ $\triangleright$ Update node values
24:  end if
25:end for
26:return $x$
@@ C.1 XRO: eXtended nonlinear Recharge Oscillator model
The XRO model is a Python implementation [89] of a recharge oscillator (RO) model for El Niño-Southern Oscillation (ENSO) [91] coupled to stochastic‐deterministic models for other climate modes (M) (as shown in Figure 11) which allows for a two-way interaction. A detailed description of the model is given by [92]. In the following we will provide an overview.
@@
The qgs***https://github.com/Climdyn/qgs/ library provides a Python implementation of a simplified, first‑order coupled atmosphere–ocean model, based in part on the Modular Arbitrary‑Order Ocean‑Atmosphere Model (MAOOAM) [93]. It represents a two‑layer quasi‑geostrophic atmosphere ($\psi^{1},\psi^{3}$), represented by barotropic $\psi_{a}$ and baroclinic $\theta_{a}$ streamfunctions, mechanically and thermally coupled to a shallow‑water ocean component ($\psi_{o}$), with interactions driven by wind forcing alongside radiative and heat exchanges (see Figure 12a).
@@
Here, $W,K,d$ and $s$ denote the coupling coefficients governing ocean–atmosphere interactions; $a,g,b$ and $c$ are the inner‑product coefficients for the atmospheric Fourier modes and $M,O,C,N$ those for the ocean. A detailed model description is given in [93] and https://qgs.readthedocs.io/en/latest/files/model/maooam_model.html.